the sales agreement that Alpha Company signed for job ABC indicated that the selling price for the job would be cost plus 150%

a. What is the price per unit using the unit product cost is $80

b. What is the price per unit using the unit product cost is $ 73.75

c. If job ABC was a test run in which Alpha company agreed to manufacture the units at cost, what is the sales price per unit?

d. If the sales agreement for the job indicated a selling price of cost plus 200%, what is the sales price per unit?
Answer #1

a)

Selling price per unit = Unit product cost + Margin

= 80 + 80 x 150%

= 80 + 120

= $200

b)

Selling price per unit = Unit product cost + Margin

= 73.75 + 73.75 x 150%

= 73.75 + 110.625

= $184.375

c)

Selling price per unit = Cost price per unit

If cost is $80, the selling price would be $80

d)

Selling price per unit = Unit product cost + Margin

= 80 + 80 x 200%

= 80 + 160

= $240